OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (akashdeepdeb):

8x - 3-5 = -8 8x -8 = -8 ADD -8 to both sides 8x -8+8 = -8+8 8x= 0 Now can you solve the whole thing!? :D

OpenStudy (anonymous):

THANKS SO MUCH!

OpenStudy (akashdeepdeb):

:)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."